# Comprehensive Benchmarking Environment for Worker Flexibility in Flexible Job Shop Scheduling Problems

### v1

@@ Abstract
In Production Scheduling, the Flexible Job Shop Scheduling Problem (FJSSP) aims to optimize a sequence of operations and assign each to an eligible machine with varying processing times. To be even closer to real production environments, the integration of the workforce may be considered. Hence, each machine also requires a worker to be present to process an operation which additionally affects the processing times. The resulting problem is called Flexible Job Shop Scheduling Problem with Worker Flexibility (FJSSP-W).
The FJSSP has been approached with different strategies considering various problem representations, including Mixed Integer Linear Programming (MILP), Constrained Programming (CP), and Simulation-based Optimization (SBO). In the latter area in particular, there exists a large number of specialized Evolutionary Algorithms (EA) like Particle Swarm Optimization (PSO) or Genetic Algorithms (GA). Yet, the solvers are often developed for single use cases only, and validated on a few selected test instances, let alone compared with results from state-of-the-art solvers using other problem representations. While suitable approaches do also exist, the design of the FJSSP-W instances is not standardized and the performance of the algorithms is hardly comparable.
This calls for a systematic benchmarking environment that provides a comprehensive set of FJSSP(-W) instances and supports targeted algorithm development. It will facilitate the comparison of algorithmic performance in the face of different problem characteristics. The present paper presents a collection of 402 commonly accepted FJSSP instances and proposes an approach to extend these with worker flexibility. The conversion of the FJSSP into an FJSSP-W has a significant impact on problem complexity. In addition, we present a detailed procedure for the evaluation of scheduling algorithms on these problem sets and provide suitable model representations for this purpose. This includes MILP formulations, CP models, and appropriate problem encodings for SBO solvers for both problem types. In addition, we provide complexity characteristics for all presented instances as well as baseline results of common commercial solvers to facilitate the validation of new algorithmic developments.
@@
[inst1]organization=Josef Ressel Centre for Robust Decision Making
Research Centre Business Informatics
Vorarlberg University of Applied Sciences,addressline=
Hochschulstrasse 1, city=Dornbirn, postcode=6850, country=Austria
@@
An alternative modeling approach for flexible job shop scheduling problems (FJSSP) as a mixed integer linear program (MILP) suitable to represent an exhaustive set of commonly accepted benchmark problems is proposed.
The paper presents a refined algorithmic approach to extend the FJSSP model with worker flexibility constraints (FJSSP-W). This way, we provide a portfolio of benchmark instances for an otherwise rather fragmented problem branch. A corresponding extension of the derived MILP representation for the introduced FJSSP-W is also delivered.
We provide a suitable problem encoding for a convenient simulation-based optimization (SBO) via meta-heuristics, e.g. Evolutionary Algorithm (EA) variants, to such problem classes and present an approach to effectively treat the worker flexibility constraints.
In addition, we develop adequate Constraint Programming (CP) representations for both problem classes (FJSSP and FJSSP-W) which facilitate the comparison with state-of-the-art solvers from this field of research.
Two problem classes in three representations (MILP, CP, and SBO) provide a comprehensive test environment for the comparison of solvers of different algorithmic paradigms. As a baseline, we compare the results of commercial state-of-the-art solvers (GUROBI, Hexaly, CPLEX (IBM), OR-Tools (Google)) and self-developed Genetic Algorithms (GA).
@@ 1 Introduction
Planning problems occur in a variety of application areas where the allocation and optimization of resources is critical to achieving specific goals. Consequently, planning problems are relevant in virtually all economic areas and the corresponding challenges are critical for ensuring sustainable economic growth, competitiveness, and efficient operations. In Supply Chain Management, for example, planning is essential for coordinating the flow of goods. In the Energy Sector, planning problems deal with the generation, distribution, and consumption of energy. Other application fields include Healthcare, Finance, Agriculture, and many more. Another important area is Manufacturing, i.e. the branch of industry that involves the production of finished products from raw materials. There, Production Scheduling (PS) deals with the allocation of resources such as machinery, labor and materials in order to achieve production targets efficiently. Such production goals include meeting customer requirements in terms of delivery times, minimizing idle time in certain production steps or responding to machine breakdowns. In this way, effective scheduling aims to promote sustainable development, drive innovation and ensure the resilience and competitiveness of manufacturing companies.
@@
Production Scheduling problems involve determining the proper sequence of actions or decisions to attain (near) optimal performance on the shop floor. They take a variety of forms and are commonly approached by applying techniques from Artificial Intelligence (AI) or Operations Research (OR). A broad overview of the PS landscape is provided in Fig. 1, in which the problem branches on which this work is based are indicated by the blue boxes.
@@
Figure 1: Localization of the problem classes under consideration in the area of Production Scheduling. Notice that other problem types exist and the illustration does by no means claim to be comprehensive.
@@
Among such tasks, the Job Shop Scheduling Problem (JSSP) is a type of scheduling problem where multiple jobs (or tasks) share a common set of resources but have different processing requirements and routes through the production facility. Each job may require different machines or processes, and the order in which the jobs are processed can affect the overall efficiency of the system. JSSPs may turn out very complex due to the variability in job processing times, resource constraints, and the need to optimize the sequence of operations. Unlike the traditional JSSP, Flexible Job Shop Scheduling Problem (FJSSP) allows jobs to be processed on different machines. This provides more flexibility in resource allocation to those machines that possesses the required capabilities for a job. In this respect, the FJSSP represents an extension of the JSSP that considers a more dynamic and adaptable environment. The flexibility allows for more dynamic and adaptive scheduling, but it also introduces additional complexity to the scheduling problem, and optimization algorithms need to consider the various combinations of machine assignments to achieve an (near) optimal schedule. An additional layer of complexity to the scheduling process is added when worker-related dynamics are also considered within a FJSSP. Such problems are commonly referred to as Flexible Job Shop Scheduling Problem with Worker Flexibility (FJSSP-W). There are many restrictions that may arise when considering worker flexibility in FJSSP, e.g., consecutive working hours, breaks or different qualifications of workers.
Depending on the modeling approach used and the optimization technique envisaged, JSSP and FJSSP model representations come in different forms. Common problem formulations include mathematical graphs, such as Resource-Task-Network (RTN) representations, permutation-based representation, or mathematical programming (MP) formulations in the form of objective functions and constraints. The latter usually come in the form of Mixed-Integer Linear Programming (MILP) problems and can be treated by various optimization techniques. Formulations as simulation-based optimization (SBO) problems are also frequently found in the literature. Instead of an analytical representation of the problem, the objectives and dependencies of a production line are implemented in the form of a black-box environment that allows the evaluation of individual production plans. Constraint Programming (CP) formulations are another way of expressing production planning problems. CP enables the integration of constraints and their solution mechanisms into a programming language. The problem is formulated in declarative form, while the solution process is secondary from the users’ point of view. This task is performed by a specialized constraint solver.
For SBO representations common approaches comprise meta-heuristic algorithms that rely on nature-inspired working principles, such as Genetic Algorithm (GA), Ant Colony Optimization (ACO) or Particle Swarm Optimization (PSO) variants. Hybrid approaches that combine multiple optimization techniques to exploit complementary strengths are also regularly considered. An example would be the combination of a GA with a Local Search method. Focusing on the exploration of the neighborhood of a given solution, Local Search (LS) solvers are used to refine solutions obtained from other methods and to find locally optimal solutions. In case of a MILP or CP formulation of the FJSSP, specialized commercial mathematical programming solvers of e.g. GUROBI and CPLEX (IBM) or constraint programming solvers like OR-Tools(Google), Hexaly or CP of CPLEX respectively, are widely used to find (near) optimal solutions. These solvers obtain state-of-the-art solutions for many FJSSP instances. Still, the choice of solver depends on factors such as the size and complexity of the FJSSP instance, the desired solution quality, and the available computational resources. Researchers and practitioners often experiment with multiple approaches to find the most effective solution for a specific FJSSP instance.
While the performance of most solvers is assessed on individual real-world problem instances and compared to selected solver variants only, some scheduling approaches are validated on rather small sets of test problem instances. In contrast to the real-world problems considered, the majority of FJSSP test sets must be considered fairly academic. Nevertheless, recent literature offers access to a number of standard FJSSP benchmark instances being spread across individual publications and modeled in different ways. Further, the benchmark problems usually vary in their structure and complexity. A unified benchmarking framework which allows dedicated algorithm performance comparisons on FJSSPs does not exist to the best of our knowledge. Moreover, there are hardly any comparable benchmark instances for FJSSP-W. As the modeling of FJSSP-W requires a careful consideration of both machine and worker-related constraints. The type of model affects the selection of solvers able to handle the increased complexity introduced by flexible scheduling policies. This introduces novel objectives like not only optimizing production efficiency but also taking into account the necessities of the workforce.
This paper proposes a systematic test environment of FJSSP and FJSSP-W instances to benchmark state-of-the-art solvers. For this purpose, an alternative MILP modeling approach for FJSSPs suitable for incorporation of a comprehensive set of 402 commonly accepted benchmark problems is introduced. Our MILP formulation allows for an improvement of the schedules found by the commercial solver GUROBI on some of the benchmark instances. The paper presents an algorithmic approach to extend the FJSSP model with worker flexibility constraints (FJSSP-W). This way, we provide a way to systematize the fragmented area of FJSSP-W instances. It provides an opportunity to create a portfolio of benchmark instances for the rather fragmented area of FJSSP-W instances. Further, suitable SBO problem encodings convenient for the application of meta-heuristics, e.g. Evolutionary Algorithm (EA) variants, as well as adequate CP representations are provided. As a result, a comprehensive test environment for the comparison of solvers from three different paradigms (MILP, CP, and SBO) on both problem classes (of 402 instances each) is established. Lastly, the paper describes a structured benchmark environment for the FJSSP(-W) which offers the possibility to choose among different problem characteristics, as well as the evaluation of a set of performance indicators (relative solution quality, wall clock time, etc.) and corresponding graphical visualizations. The test environment also provides characteristic attributes for each problem instance, such as the total number of operations, the flexibility of the machines or the problem-related duration variety, on the basis of which a targeted selection of sub-instances can be made. The functionalities are demonstrated by presenting baseline results of various state-of-the-art commercial solvers (GUROBI, Local Solver, CPLEX (IBM), OR-Tools (Google)) and a self-developed Genetic Algorithms (GA).
The remainder of this paper is structured as follows: In Section 2, a detailed description of the considered problem classes is provided. In addition, related work is discussed and available benchmarks as well as state-of-the-art solvers are presented. Section 3 deals with the formulation of the respective FJSSP representations. Next, the procedure for integration of worker flexibility is specified in Sec. 4. It realizes the transformation of an FJSSP instance into an FJSSP-W instance with just a few user-defined assumptions. Subsequently, the corresponding FJSSP-W formulations are specified. The detailed description of the benchmarking environment and its configuration options are the subject of Sec. 5. Afterwards, Section 6 presents a number of baseline results obtained by application of several state-of-the-art solvers. The contribution of this work and its impact on the field of research is discussed in summary in Sec. 7. Finally, Section 8 gives an outlook on future extensions of the presented benchmarking suite.
@@ 2 Problem Formulation & Related Work
The Flexible Job Shop Scheduling Problem (FJSSP) is the starting point for the following investigations. In this section, a formal definition of this problem is presented in Sec. 2.1 and current work on the FJSSP is outlined in Sec. 2.3. A large proportion of such research looks at very individual problems from real-world applications or only compares the developed solvers on individually selected benchmark instances. This complicates the performance comparison between the solvers significantly. In order to counteract this situation and to develop a structured benchmarking environment, Sec. 2.2 provides an overview of the available benchmark instances in the field of FJSSP and their characteristics.
@@ 2.1 Problem Formulation
In general formulation, the FJSSP introduces a set of $n$ jobs $\mathcal{J}=\{J_{1},\dots,J_{n}\}$ that need to be processed on a set of $m$ machines =$\mathcal{M}=\{M_{1},\dots,M_{m}\}$. Each of these jobs $J_{i}\>(i=1,\dots,n)$ consists of a predetermined sequence of $n_{i}$ operations. To this end, the notation $O_{i,j}$ refers to the $j$th operation within job $i$. Each of these operations $O_{i,j}$ is performed by one of the admissible machines in the set $\mathcal{M}_{i,j}\subset\mathcal{M}$. If operation $O_{i,j}$ is assigned to a specific machine, this machine is denoted by $M_{i,j}$. The total number of operations in a FJSSP can thus be calculated as $\texttt{N}=\sum_{i=1}^{n}n_{i}$. Accordingly, the ratio of the total amount of operations and the number of jobs gives the average number of operations per job $\frac{\texttt{N}}{n}$. The constant processing time of operation $O_{i,j}$ on machine $M_{k}\in\mathcal{M}_{i,j}\>(k=1,\dots,m)$ is given as $T_{i,j,k}$. The minimum processing duration among all possible operation-machine combinations of a single problem instance is referred to as ${T}_{\min}$. Conversely, the maximum processing duration is ${T}_{\max}$. Their difference defines the so-called duration span $\Delta T={T}_{\max}-{T}_{\min}$. In common scenarios, the general representation of an FJSSP assumes that all orders are known and all machines are operational at time 00.
In the context of FJSSPs, a large number of objective functions are conceivable with respect to the optimization of which the appropriate schedule must be composed. The performance criteria include minimizing the makespan, minimizing tardiness or maximizing resource utilization, to name but a few. While the makespan, i.e. the total time required for the execution of all jobs of a problem instance, is one of the most important objective functions in literature and it can be used as an illustrative example in the remainder of this paper, the objective function is to be represented as generally as possible. This allows the subsequent inclusion of various optimization objectives both in the problem modeling and in the performance analysis within the proposed benchmarking environment. Since every maximization task can be formulated equivalently as a minimization problem, w.l.o.g. it can be assumed that a given cost function is to be minimized in full satisfaction of the production constraints of the FJSSP formulated above.
@@
As it is common in the field of meta-heuristic optimization, the expression fitness will be used synonymously for the quality of a discovered feasible solution $\mathbf{y}$, i.e. a feasible schedule, in the subsequent sections of this work. That is, solution candidates of higher fitness (or quality) correspond to solutions of lower cost function values. The quality of two admissible scheduling solutions can consequently be compared on the basis of their objective function values. Accordingly, lower cost function values $C(\mathbf{y})$ correspond to better production plans $\mathbf{y}\in\mathcal{Y}$. For a proper comparison of two specific schedules, an order relation $\prec_{c}$ is defined by
@@
Hence, the fitness of $\mathbf{y}$ would be rated higher than that of $\mathbf{z}$ in the case of Eq. (1). The corresponding smaller or equal relation $\preceq_{c}$ is defined analogously.
2.2 Existing FJSSP Benchmarks
This section discusses the existing benchmark instances for the Flexible Job Shop Scheduling Problem. While there are already a number of test problem collections available for the FJSSP, these are often not systematically used for the evaluation of the developed solvers. Thus, the number of algorithms that are benchmarked on a large number of different FJSSP instances is relatively small [2]. A majority of solvers are rather evaluated on problem subgroups originating from distinct sources, which means that the results of such studies are always subject to a latent problem selection bias. A large number of papers also only consider individual real-world problem representations. The algorithms developed in this way are highly specialized and are rarely compared carefully with established solvers. This has two potential implications: on the one hand, the advantages of rather good solvers may remain unclear if only demonstrated in a very specific setting and secondly, incremental approaches may be overvalued if they are not compared with state-of-the-art methods.

### v2

# A Benchmarking Environment for Worker Flexibility in Flexible Job Shop Scheduling Problems

@@ Abstract
In Production Scheduling, the Flexible Job Shop Scheduling Problem (FJSSP) aims to optimize a sequence of operations and assign each to an eligible machine with varying processing times. To be even closer to real production environments, the integration of the workforce may be considered. Hence, each machine also requires a worker to be present to process an operation which additionally affects the processing times. The resulting problem is called Flexible Job Shop Scheduling Problem with Worker Flexibility. Such problems are approached with different strategies considering various problem representations, including Mixed Integer Linear Programming, Constrained Programming, and Simulation-based Optimization. In the latter area, in particular, there exists a large number of specialized Evolutionary Algorithms. Yet, solvers are often developed for distinct use cases only, and validated on few selected test instances, let alone compared with results from multiple state-of-the-art solvers. While suitable approaches exist, the design of the FJSSP-W instances is not standardized and algorithm performance is hardly comparable. This requires a systematic benchmarking environment that provides a comprehensive set of FJSSP(-W) instances for targeted algorithm development. It will facilitate comparing algorithmic performance in the face of different problem characteristics. The present paper presents a collection of 402 commonly accepted FJSSP instances and proposes an approach to extend these with worker flexibility. Additionally, it presents an environment for the evaluation of scheduling algorithms on these problem sets. Further, instance characteristics and baseline results of common commercial solvers are presented to facilitate the validation of new algorithmic developments.
@@
[inst1]organization=Research Centre Business Informatics, Josef Ressel Centre for Robust Decision Making,
Vorarlberg University of Applied Sciences,addressline=Hochschulstrasse 1, city=Dornbirn, postcode=6850, country=Austria
@@
Introduces an algorithmic method to enhance FJSSP with worker flexibility constraints.
Establishes a portfolio of benchmark instances for FJSSP with Worker Flexibility (FJSSP-W).
Presents an open-source benchmarking environment for FJSSP and FJSSP-W solvers.
Supports algorithm development with user-specific filtering and harmonized results.
Provides problem encoding for multiple optimization paradigms: meta-heuristics, MILP, and CP.
@@ 1 Introduction
Planning problems arise in various fields where resource allocation and optimization are crucial for achieving goals. These problems are relevant across all economic sectors, impacting sustainable growth, competitiveness, and efficiency [1, 2, 3].
@@
In Supply Chain Management, planning coordinates the flow of goods. In the Energy Sector, it addresses the generation, distribution, and consumption of energy. Other fields include Healthcare, Finance, and Agriculture. In Manufacturing, Production Scheduling (PS) allocates resources like machinery, labor, and materials to meet production targets efficiently. Goals include meeting delivery times, minimizing idle time, and responding to machine breakdowns. Effective scheduling promotes sustainable development, drives innovation, and ensures the resilience and competitiveness of manufacturing companies.
@@
Production Scheduling problems involve determining the proper sequence of actions or decisions to attain (near) optimal performance on the shop floor. They take a variety of forms and are commonly approached by applying techniques from Artificial Intelligence (AI), Operations Research (OR) or Simulation-based Optimization (SBO) [2]. A broad overview of the PS landscape is provided in Fig. 1, in which the problem branches on which this work is based are indicated by the blue boxes.
@@
Figure 1: Visualization of the problem classes in the area of Production Scheduling considered in this paper. Notice that other problem types exist and the illustration does by no means claim to be comprehensive.
@@
The Job Shop Scheduling Problem (JSSP) involves scheduling multiple jobs that share resources but have different processing requirements and routes. Each job may need different machines or processes, and the order of processing affects system efficiency. JSSPs are complex due to variable job processing times, resource constraints, and the need to optimize operation sequences. Unlike the classical JSSP, Flexible Job Shop Scheduling Problem (FJSSP) allow jobs to be processed on different machines. This provides more flexibility in resource allocation and may be regarded an extension of the JSSP that considers a more dynamic and adaptable environment. Yet, the flexibility introduces additional complexity, as optimization algorithms need to deal with the combinations of machine assignments to achieve an (near) optimal schedule. In case that workforce-related dependencies are included, the problem is known as Flexible Job Shop Scheduling Problem with Worker Flexibility (FJSSP-W). This adds complexity due to factors like consecutive working hours, breaks, and worker qualifications on the individual machines.
A comprehensive collection and descriptions of commonly used FJSSP instances are provided in [4]. The paper also proposes some additional FJSSP benchmark instances and provides best-known results for all reported instances. In 2024, a more recent survey [3] discusses common problem instances, different problem constraints, as well as solution approaches. While the paper updates the best-known results for the discussed FJSSP instances with more recent findings, it must be noted that there are problem instances in both works that are not considered in the respective other work. The paper concludes that the most important optimization criterion is the time span [3], although other criteria are also discussed. These include machine utilization, delays, weighted lead times, energy costs, queue times and just-in-time measures. Additional constraints such as minimum waiting times between two operations, a maximum waiting time for starting orders in the schedule, resource availability constraints, batch formation constraints, setup time constraints, resource blocking constraints (e.g. limited queue space at the machines), transportation constraints and some other special cases can be found in the literature, but they do not occur to a significant extent.
As far as the systematic benchmarking [5, 6] of algorithmic solvers for such PS problems is concerned, only a few partial aspects are covered, particularly in the area of SBO methods. The development and the performance comparison of of such solvers particularly relies on benchmarking due to the lack of theoretical performance results for optimization tasks of notable complexity. First and foremost, benchmarking experiments are established for performance evaluation and algorithm comparison on given problem classes. Ideally, this is supposed to support the selection of the algorithm best suitable for given real-world applications [7]. Further, benchmarks can be used to deliver insights into the working principles of algorithms and foster their development for specific problem types [8, 9].
Unlike other problem areas such as continuous single objective constrained optimization problems [10] or combinatorial Traveling Sales Person problems [11], to the best of our knowledge, there is no commonly accepted benchmarking environment for the FJSSP (Flexible Job Shop Scheduling Problem) for the aforementioned purposes. Instead, the situation is segmented into the following categories:
Sets of problem instances
Over the years, multiple authors proposed their own versions of FJSSP test problem instance sets [4, 12]. These problem sets are widely accepted and used for the benchmarking of algorithmic developments. In addition, new problem instances are constantly being created [13]. Although, comprehensive collections of such sets are openly available [14, 15, 16, 4], the rather arbitrary use of instance subsets is the usual practice. This complicates the cross-comparability of solvers and impedes a preferably complete picture of the situation.
Real-world problem instances
There are many individual real-world problems that are used to justify algorithmic developments. While there are efforts to derive reasonable benchmarking environments form specific industrial use cases [17, 18], the majority of cases does not allow for the (systematically) comparison of different solvers. This is partly due to the real-world problems coming with very individual requirements that complicate the application of many solvers [19]. On the other hand, the computational effort can be too extensive for a rigorous assessment. Moreover, not all details of the real-world problem dependencies might be properly disclosed in all cases [18].
Competitions
In the context of conferences such as GECCO [^1] An example is the GECCO2025 Competition on Evolutionary Computation in the Energy Domain: Summer Finals of the Risk-based Energy Scheduling which is held during the Genetic and Evolutionary Computation Conference in 2025 https://www.gecad.isep.ipp.pt/ERM-competitions/2025-2/. and CEC [^2] For example, in the context of the Congress on Evolutionary Computation, the CEC’2024 Competition on Large-Scale Sequence Optimization for Production Lines was organized http://developer.xuelangyun.com/#/competition?competitionId=1009000&lang=en. In 2025, the same conference hosts the CEC’2025 Competition on Dynamic Scheduling of Flexible Flow Shop with Random Job Arrivalshttps://github.com/SIA-aps/CEC2025-DFFSP-competition. , competitions on scheduling problems are regularly announced. These are usually well documented and structured so that they can also be used for benchmarking purposes. However, these competitions are usually also tailored to specific use cases from the areas of production or planning, sometimes with very special requirements. A broader comparison with solvers that have not been adapted to these problem-specific particularities is therefore difficult.
Performance metrics
A frequently discussed benchmarking aspect in Scheduling Problems is the definition and selection of performance metrics [20, 21]. These can vary greatly depending on the scheduling objective, application area and other influencing factors (noise, dynamic constraints, etc.) [22]. While it is hardly possible to prescribe a uniform guideline in this respect, it is important to pay close attention to the consistent use of such performance measures when comparing algorithms.
When it comes to considering worker-related constraints, i.e. the FJSSP-W, the benchmarking landscape becomes even more scattered. There is neither a commonly accepted set of FJSSP-W instances nor a standardized method for creating FJSSP-W test instances. Instead, researchers often develop test problem instances independently [23, 24] or rely on solitary problem representations from the real world [25, 26]. The lack of standardization complicates benchmarking principles such as comparability and reproducibility, as documentation may be incomplete or not all aspects of the real-world problem are disclosed.
Consequently, this paper presents a benchmark environment that summarizes a comprehensive collection of problem instances for both the FJSSP as well as the FJSSP-W and allows for a systematic assessment of corresponding solvers. To this end, those FJSSP instances from the collections of [4] and [3] are selected that fulfill the FJSSP definition and must not be assigned to more specialized problem classes (i.e. parallel machines, or flow store scheduling problems, cf. Fig. 1). Such FJSSP instances are then extended to FJSSP-W instances. Through the traceable modification, the increase in complexity between the problem classes and its effects on algorithm performance can be better understood.
The benchmark environment represents an open-access platform for the performance evaluation and comparison of scheduling algorithms for the problem classes mentioned. By specifying standardized test criteria and termination conditions, benchmarking can be guaranteed to be as harmonized as possible. This is accompanied by a coherent visualization of the benchmarking results. The platform will affect the relative performance assessments of state-of-the-art solvers and be useful for the selection of solvers for specific problem classes. Additionally, it supports the development of algorithms for specific problem characteristics, such as particularly high flexibilities in the production environment, through optional user-specific filtering based on the characteristics of the provided benchmark instances. By selecting a smaller number of representative instances for a particular problem property, algorithmic developments can be tested repeatedly with less computational effort before they have to be applied to the entirety of the instances.
The benchmarking environment is designed primarily for the comparison and development of simulation-based optimization approaches, e.g. Evolutionary Algorithms (EA), and is nevertheless intended to be accessible for classical Mathematical Programming methods. To this end, the supplementary material of this paper includes Mixed-Integer Linear Programming (MILP) and Constraint Programming (CP) problem representations. These models also allow the reproducibility of the best-known results available in the benchmarking environment, which were obtained by applying multiple state-of-the-art solvers like Gurobi, Hexaly, CPLEX, or OR-Tools [27, 28, 29, 30].
The paper is organized as follows: Section 2 elaborates on the FJSSP formulation and provides an overview of available benchmark sets. The extension towards FJSSP-W is then specified in Sec.3. In Sec.4, the benchmarking environment is introduced. This is followed by the demonstration of sample results of several test solvers in Sec.5. Finally, the paper concludes in Sec.6, and proposes some directions for future work in Sec.7, respectively.
@@ 2 Problem Formulation & Related Work
The Flexible Job Shop Scheduling Problem (FJSSP) is the starting point for the following investigations. In this section, a formal definition of this problem is presented in Sec. 2.1 together with the introduction of necessary problem features (i.e. flexibility or duration variety) that allow for the characterization of individual problem instances. Further, Sec. 2.2 provides an overview of the available benchmark instances in the field of FJSSP.
@@ 2.1 Problem Formulation
Using the notation from [32], the general representation of a FJSSP introduces a set of $n$ jobs $\mathcal{J}=\{J_{1},\dots,J_{n}\}$ that need to be processed on a set of $m$ machines =$\mathcal{M}=\{M_{1},\dots,M_{m}\}$. Each of these jobs $J_{i}\>(i=1,\dots,n)$ consists of a predetermined sequence of $n_{i}$ operations. To this end, the notation $O_{i,j}$ refers to the $j$th operation within job $i$. Each of these operations $O_{i,j}$ is performed by one of the $m_{i,j}$ admissible machines in the set $\mathcal{M}_{i,j}\subset\mathcal{M}$. If operation $O_{i,j}$ is assigned to a specific machine, this machine is denoted by $M_{i,j}$. The total number of operations in a FJSSP can thus be calculated as $\texttt{N}=\sum_{i=1}^{n}n_{i}$. Accordingly, the ratio of the total amount of operations and the number of jobs gives the average number of operations per job $\frac{\texttt{N}}{n}$. The constant processing time of operation $O_{i,j}$ on machine $M_{k}\in\mathcal{M}_{i,j}\>(k=1,\dots,m)$ is given as $T_{i,j,k}$. The minimum processing duration among all possible operation-machine combinations of a single problem instance is referred to as ${T}_{\min}$. Conversely, the maximum processing duration is ${T}_{\max}$. Their difference defines the so-called duration span $\Delta T={T}_{\max}-{T}_{\min}$. In common scenarios, the general representation of an FJSSP assumes that all orders are known and all machines are operational at time 00.
In the context of FJSSPs, many objective functions are conceivable, the optimization of which determines the final production plan. The performance criteria include minimizing the makespan, minimizing tardiness or maximizing resource utilization, to name but a few. While the makespan, i.e. the total time required for the execution of all jobs of a problem instance, is one of the most important objective functions in literature and it can be used as an illustrative example in the remainder of this paper, the objective function is to be represented as generally as possible. This allows the subsequent inclusion of various optimization objectives both in the problem modeling and in the performance analysis within the proposed benchmarking environment. Since every maximization task can be formulated equivalently as a minimization problem, w.l.o.g. it can be assumed that a given cost function is to be minimized in full satisfaction of the production constraints of the FJSSP formulated above.
@@
As it is common in the field of meta-heuristic optimization, the term fitness will be used synonymously for the quality of a discovered feasible solution $\mathbf{y}$, i.e. a feasible schedule, in the subsequent sections of this work. Hence, candidate solutions of higher fitness (or quality) correspond to solutions of lower cost function values. The quality of two admissible scheduling solutions can consequently be compared on the basis of their objective function values. Accordingly, lower cost function values $C(\mathbf{y})$ correspond to better production plans $\mathbf{y}\in\mathcal{Y}$. For a the comparison of two feasible schedules $\mathbf{y}$ and $\mathbf{z}$, an order relation $\prec_{c}$ is defined by
@@
That is, in this case the fitness of $\mathbf{y}$ is rated higher than that of $\mathbf{z}$ with respect to the cost function $C$ (e.g. the makespan of the schedules). The corresponding smaller or equal relation $\preceq_{c}$ is defined analogously.
2.2 Foundational FJSSP benchmark problem instances
This section discusses the existing benchmark instances for the Flexible Job Shop Scheduling Problem. While there are already a number of test problem collections available for the FJSSP, these are often not systematically used for the evaluation of the developed solvers. Thus, the number of algorithms that are benchmarked on a large number of different FJSSP instances is relatively small [3]. A majority of solvers are rather evaluated on problem subgroups originating from distinct sources, which means that the results of such studies are always subject to a latent problem selection bias. A large number of papers also only consider individual real-world problem representations [2]. The algorithms developed in this way are highly specialized and are rarely compared carefully with established solvers. This has two potential implications: on the one hand, the advantages of rather good solvers may remain unclear if only demonstrated in a very specific setting and secondly, incremental approaches may be overvalued if they are not compared with state-of-the-art methods.